OpenStudy (anonymous):

Use Property 2 for radicals to simplify the following radical expression as much as possible. Assume all variables represent positive numbers.

OpenStudy (anonymous):

\[\sqrt[3]{\frac{ 64m^3 }{ 27n^3 }}\]

OpenStudy (anonymous):

@kx2bay

OpenStudy (anonymous):

what's the cubic root of 64 , also of 27 ?

OpenStudy (anonymous):

how much is 3*3*3 and 4*4*4

OpenStudy (anonymous):

cookie, you watching family guy again?

OpenStudy (anonymous):

meant cookii :)

OpenStudy (anonymous):

lol, no, my friend called me lol i wish i was

OpenStudy (anonymous):

ahhh i felt like you were absent for a bit

OpenStudy (anonymous):

I'm back :)

OpenStudy (anonymous):



OpenStudy (anonymous):

is the above correct implementation of prop.2 ? know how to simplify it further? I was asked to make you think and not give you the answer straight away, so that why this torture ;)

OpenStudy (anonymous):

lol who put you up to this?

OpenStudy (anonymous):

3 &4

OpenStudy (anonymous):

x^3=x^2*x

OpenStudy (anonymous):

the moderator yes 3 and 4 are the answers to the number, so now can you simplify the the rest ?

OpenStudy (anonymous):

know the answer, or still struggling ?

OpenStudy (anonymous):

uhh, I think I do, just donw how to set it up lol

OpenStudy (anonymous):

4/3

OpenStudy (anonymous):

does it look like this ;) 4m / 3n

OpenStudy (anonymous):

that sentence was all messed up lol

OpenStudy (anonymous):

it do

OpenStudy (anonymous):

ok, dont tell the moderators that i didn't make u think, ok ??? lol

OpenStudy (anonymous):

i think you're even sweating from too much thnking, right ;)

OpenStudy (anonymous):

lol i'm sweating period...California is hot ! lol

OpenStudy (anonymous):

hahaha...its bloody cold here, got the heater next to me

OpenStudy (anonymous):

virtually sitting on it ;)

OpenStudy (anonymous):

My fan is on it's highest lol